If the sum of all edges of cube is 36 then find its volume
Found 3 solutions by TimothyLamb, Alan3354, Edwin McCravy:
Answer by TimothyLamb(4379)

a cube has 12 edges, all of equal length:
---
36/12 = 3
---
v = 3*3*3
v = 27
---
Free algebra tutoring live chat:
https://sooeet.com/chat.php?gn=algebra
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations with quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php
---

Answer by Alan3354(69443)

If the sum of all edges of cube is 36 then find its volume
---------
A cube has 8 equal edges of length s.
8s = 36
s = 4.5
----
Vol = s^3 = 4.5^3
Vol = 91.125 cubic units

Answer by Edwin McCravy(20081)

A cube has 4 edges on the front square face, 4 edges on the back square face,
and 4 edges connecting the corners of the front square face to the
corners of the back square face.





That's 12 faces. So since the sum of all 12 edges is 36, each edge is 36/12 
or 3.  The volume of a cube is the length of an edge raised to the third
power. [That's why they call raising to the third power "CUBING"].

Volume = (edge)3 = 33 = 27.

Look at the picture, There are 9 1x1x1 cubes in the front, 9 in the middle,
and 9 in the back.  That's a volume of 27 small 1x1 cubes.

Edwin